NOAA GOES SATOPS Morning Report: Tuesday, February 08, 2005

Status Report From: National Oceanic and Atmospheric Administration
Posted: Tuesday, February 8, 2005

DAY 39

NOTE: * GOES-12 Maneuver schedule for today (DOY039) at 1718Z.


OPS 48 rev 40511 at POGO-B 07:39 Marginal Noisy 2K telemetry, data delayed but good to user. MIR submitted.

OPS 47 rev 50983 at POGO-B 07:44 Failed Noisy telemetry on the CH-8 2K only, the data on CH-2 was delayed but good and

POGO-B's trouble was a bit sync which was swapped out for spare unit, the next support was good. Clockset not done. MIR submitted.

All other DMSP operations were nominal over the past 24 hours.


N15 rev 35032 at Fairbanks 1655Z: Lost HRPT TLM from 17:00:18 to 17:02:05 (1 minute 47 seconds), due to bit sync problem on string #1. G4B short: 94% recovered.

N15 rev 35039 at Fairbanks 0429Z: All EPOCH & DOWNLINK directives failed - FCDA manually establish command capability. Unable to accomplish post-pass transfers from 13m 'C' IMUX, due to FEP problems, but 13m 'B' shadowed the pass, so transfers were recovered from it. FEP software was recycled after the pass (and the problem) was corrected.

All other POES operations were nominal over the past 24 hours.


All GOES operations were nominal over the past 24 hours.

Questions or comments to

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter